# T-FIX: Text-Based Explanations with Features Interpretable to eXperts

### v1

@@ 9 Conclusion
GPT-4o
Vanilla 0.039 0.103 0.916 0.259 0.075 0.567 0.657
Chain-of-Thought 0.044 0.093 0.824 0.286 0.103 0.460 0.714
Socratic Prompting 0.044 0.127 0.829 0.277 0.115 0.462 0.657
SubQ Decomposition 0.049 0.118 0.837 0.304 0.115 0.485 0.657
o1
Vanilla 0.044 0.170 0.784 0.304 0.194 0.656 0.752
Chain-of-Thought 0.045 0.146 0.818 0.339 0.177 0.685 0.750
Socratic Prompting 0.042 0.155 0.793 0.348 0.155 0.646 0.755
SubQ Decomposition 0.044 0.147 0.818 0.321 0.138 0.695 0.780
Gemini-2.0-Flash
Vanilla 0.045 0.145 0.831 0.223 0.253 0.577 0.654
Chain-of-Thought 0.042 0.118 0.837 0.232 0.255 0.558 0.663
Socratic Prompting 0.041 0.118 0.809 0.232 0.159 0.592 0.661
SubQ Decomposition 0.053 0.109 0.773 0.241 0.249 0.562 0.688
Claude-3.5-Sonnet
Vanilla 0.053 0.127 0.962 0.241 0.146 0.485 0.709
Chain-of-Thought 0.050 0.118 1.012 0.268 0.150 0.538 0.735
Socratic Prompting 0.044 0.118 0.998 0.232 0.145 0.508 0.748
SubQ Decomposition 0.050 0.136 0.990 0.259 0.149 0.485 0.741
Table A2: Evaluating top LLMs on T-FIX. We report the average performance of the LLM across all examples in the dataset. We report accuracy for classification tasks, and MSE for regression tasks – a () indicates that the score reported is MSE. Baseline implementations are described in Section 6.
@@
Appendix A Extending T-FIX to a New Domain
To add a new domain to T-FIX, we advise you to follow these steps:
  1. 1.
    Generate criteria: Use the deep research prompt template shown in Figure˜A4 to generate a list of expert alignment criteria for your domain. Optionally, have a domain expert vet the generated criteria.
  2. 2.
    Modify prompts: Modify the prompt templates outlined in Figure˜A1, Figure˜A2, and Figure˜A3 with your task description, few-shot examples, and generated expert criteria.
  3. 3.
Appendix B Prompts for T-FIX Pipeline
We show the prompts for Stage 1, 2, and 3 in Figure˜A1, Figure˜A2, and Figure˜A3, respectively. These prompts show a high-level template that was used by all domains. In practice, authors iterated multiple times on each domain’s prompts, experimenting with the instruction wording and few-shot examples that yielded the best possible results.
@@ Appendix C Ablation for Evaluation Models
To assess the robustness of our evaluation pipeline, we conduct ablations using different models as evaluators: closed-sourced GPT-4o (openai2024gpt4ocard), Gemini-2.0-Flash (google2024gemini2flash), and an open-source model Qwen2.5-VL-7B-Instruct (Qwen2.5-VL). We evaluate vanilla prompts on three datasets, each representing a different domain: Mass Maps (cosmology), Emotion (psychology), and Cholecystectomy (medicine).
Table˜A1 summarizes the results:
  • Mass Maps (cosmology): All three evaluators unanimously rank Claude-3.5-Sonnet as the best model and o1 as the second best.
  • Cholecystectomy (medicine): GPT-4o and Qwen2.5-VL-7B-Instruct both rank o1 as the best and Qwen2.5-VL-7B-Instruct as the second best. Gemini-2.0-Flash is more conservative, labeling all claims as irrelevant and assigning all models a score of 0.
  • Emotion (psychology): All evaluators assign similar scores across generation models, leading to minimal differences in rankings.
@@
Figure A2: Prompt Template for Stage 2: Relevancy Filtering
@@
Appendix D T-FIX Datasets: Additional Details
@@ D.1 Mass Maps
Data Selection & Preprocessing.
We randomly sampled 100 examples from the MassMaps test set. To ensure compatibility with LLMs like GPT-4o, which operate on a 32×32 patch size, we upsampled each image by a factor of 11 to preserve spatial detail and avoid patch-level compression. Instead of raw pixel values, we applied a colormap based on expert-defined intensity thresholds used to identify key cosmological features such as voids and clusters. Pixel intensities were scaled by standard deviations to emphasize meaningful variation. We found that larger, visually enhanced inputs reduced refusal rates from LLMs and encouraged more consistent responses.
Explanation Prompt.
Figure A6 shows the prompt used to generate LLM explanations for predicting $\Omega_{m}$ and $\sigma_{8}$. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5. The prompt includes a description of how pixel values are mapped to colors, as well as the valid ranges for $\Omega_{m}$ and $\sigma_{8}$. Without this range, models tend to default to common values (e.g., 0.3 for $\Omega_{m}$, 0.8 for $\sigma_{8}$), reducing response variability.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:
@@
D.2 Supernova
Task.
The objective is to classify astrophysical objects using time-series data comprising observation times (Modified Julian Dates), wavelengths (filters), flux values, and corresponding flux uncertainties. We use data from the PLAsTiCC challenge (theplasticcteam2018photometric), where the model must predict one of 14 astrophysical classes.
Data Selection & Preprocessing.
We sampled 100 examples across the Supernova train, validation, and test sets, aiming for 7–8 instances per class to mitigate class imbalance. For rare classes with only one test set instance, we included all available examples from the validation and test sets, supplementing with training samples to meet the target count. For LLM input, we converted each raw time series into a multivariate time-series plot: time is on the x-axis, flux on the y-axis, error bars denote flux uncertainty, and point colors indicate different wavelengths.
Explanation Prompt.
Figure A7 shows the prompt used to generate explanations for classifying astronomical objects. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A5. The prompt includes a description of the input plot as a multivariate time series and provides the full list of possible class labels to guide the model’s predictions.
Expert Criteria.
The expert-validated criteria for expert alignment calculation are listed below:

### v2

@@ Appendix B Annotation Study Setup
GPT-5.2-Pro (OpenAI, 2025a)
Vanilla $0.060^{*}_{\pm 0.01}$ $0.236_{\pm 0.04}$ $0.801^{*}_{\pm 0.11}$ $320_{\pm 0.05} 0.0$$014^{*}_{\pm 0.01}$ $0.612_{\pm 0.04}$ $0.627_{\pm 0.05}$
Chain-of-Thought $0.058^{*}_{\pm 0.01}$ $0.218_{\pm 0.04}$ $0.868^{*}_{\pm 0.12}$ $0.340_{\pm 0.05}$ $0.162^{*}_{\pm 0.02}$ $0.623_{\pm 0.04}$ $0.618_{\pm 0.05}$
Socratic Prompting $0.062^{*}_{\pm 0.01}$ $0.218_{\pm 0.04}$ $0.841^{*}_{\pm 0.12}$ $300_{\pm 0.05} 0.0$$098^{*}_{\pm 0.02}$ $0.623_{\pm 0.04}$ $0.637_{\pm 0.05}$
SubQ Decomposition $0.061^{*}_{\pm 0.01}$ $0.227_{\pm 0.04}$ $0.871^{*}_{\pm 0.12}$ $0.340_{\pm 0.05}$ $0.146^{*}_{\pm 0.02}$ $0.623_{\pm 0.04}$ $0.627_{\pm 0.05}$
GPT-5-mini (OpenAI, 2025b)
Vanilla $0.048^{*}_{\pm 0.00}$ $0.191_{\pm 0.04}$ $1.004^{*}_{\pm 0.15}$ $0.410_{\pm 0.05}$ $0.149^{*}_{\pm 0.01}$ $0.585_{\pm 0.04}$ $0.624_{\pm 0.05}$
Chain-of-Thought $0.046^{*}_{\pm 0.00}$ $0.164_{\pm 0.04}$ $1.048^{*}_{\pm 0.14}$ $350_{\pm 0.05} 0.0$$067^{*}_{\pm 0.01}$ $0.577_{\pm 0.04}$ $0.676_{\pm 0.05}$
Socratic Prompting $0.047^{*}_{\pm 0.00}$ $0.200_{\pm 0.04}$ $1.018^{*}_{\pm 0.15}$ $310_{\pm 0.05} 0.0$$018^{*}_{\pm 0.01}$ $0.585_{\pm 0.04}$ $0.627_{\pm 0.05}$
SubQ Decomposition $0.050^{*}_{\pm 0.00}$ $0.164_{\pm 0.03}$ $1.024^{*}_{\pm 0.15}$ $340_{\pm 0.05} 0.0$$011^{*}_{\pm 0.00}$ $0.577_{\pm 0.04}$ $0.608_{\pm 0.05}$
Claude-Opus-4.5 (Anthropic, 2025b)
Vanilla $035^{*}_{\pm 0.00} 0.0$$082_{\pm 0.03}$ $0.701^{*}_{\pm 0.08}$ $0.300_{\pm 0.05}$ $0.273^{*}_{\pm 0.01}$ $0.454_{\pm 0.05}$ $0.618_{\pm 0.05}$
Chain-of-Thought $0.036^{*}_{\pm 0.00}$ $0.118_{\pm 0.03}$ $0.814^{*}_{\pm 0.12}$ $0.300_{\pm 0.05}$ $0.287^{*}_{\pm 0.01}$ $0.469_{\pm 0.04}$ $0.634_{\pm 0.05}$
Socratic Prompting $0.035^{*}_{\pm 0.00}$ $0.109_{\pm 0.03}$ $0.708^{*}_{\pm 0.08}$ $0.310_{\pm 0.05}$ $0.261^{*}_{\pm 0.02}$ $0.462_{\pm 0.04}$ $0.620_{\pm 0.05}$
SubQ Decomposition $035^{*}_{\pm 0.00} 0.0$$082_{\pm 0.03}$ $0.774^{*}_{\pm 0.08}$ $0.300_{\pm 0.05}$ $0.277^{*}_{\pm 0.01}$ $0.469_{\pm 0.04}$ $0.618_{\pm 0.05}$
Claude-Haiku-4.5 (Anthropic, 2025a)
Vanilla $040^{*}_{\pm 0.00} 0.0$$036_{\pm 0.02}$ $0.681^{*}_{\pm 0.11}$ $0.270_{\pm 0.04}$ $0.290^{*}_{\pm 0.01}$ $0.677_{\pm 0.04}$ $0.647_{\pm 0.05}$
Chain-of-Thought $039^{*}_{\pm 0.00} 0.0$$036_{\pm 0.02}$ $0.748^{*}_{\pm 0.12}$ $0.260_{\pm 0.04}$ $0.284^{*}_{\pm 0.01}$ $0.669_{\pm 0.04}$ $0.608_{\pm 0.05}$
Socratic Prompting $049^{*}_{\pm 0.00} 0.0$$091_{\pm 0.03}$ $0.671^{*}_{\pm 0.10}$ $0.270_{\pm 0.05}$ $0.303^{*}_{\pm 0.01}$ $0.677_{\pm 0.04}$ $0.324_{\pm 0.05}$
SubQ Decomposition $043^{*}_{\pm 0.00} 0.0$$055_{\pm 0.02}$ $0.728^{*}_{\pm 0.11}$ $0.240_{\pm 0.04}$ $0.296^{*}_{\pm 0.01}$ $0.677_{\pm 0.04}$ $0.412_{\pm 0.05}$
Gemini-2.5-Pro (team, 2025)
Vanilla $0.097^{*}_{\pm 0.01}$ $0.227_{\pm 0.04}$ $0.901^{*}_{\pm 0.14}$ $0.280_{\pm 0.04}$ $0.216^{*}_{\pm 0.01}$ $0.408_{\pm 0.04}$ $0.598_{\pm 0.05}$
Chain-of-Thought $0.101^{*}_{\pm 0.01}$ $0.209_{\pm 0.04}$ $0.971^{*}_{\pm 0.15}$ $0.270_{\pm 0.04}$ $0.219^{*}_{\pm 0.01}$ $0.392_{\pm 0.04}$ $0.588_{\pm 0.05}$
Socratic Prompting $0.093^{*}_{\pm 0.01}$ $0.173_{\pm 0.04}$ $0.858^{*}_{\pm 0.15}$ $0.290_{\pm 0.05}$ $0.195^{*}_{\pm 0.02}$ $0.415_{\pm 0.04}$ $0.569_{\pm 0.05}$
SubQ Decomposition $0.087^{*}_{\pm 0.01}$ $0.155_{\pm 0.04}$ $0.815^{*}_{\pm 0.15}$ $0.273_{\pm 0.04}$ $0.151^{*}_{\pm 0.01}$ $0.408_{\pm 0.04}$ $0.608_{\pm 0.05}$
Gemini-2.5-Flash (team, 2025)
Vanilla $081^{*}_{\pm 0.01} 0.0$$082_{\pm 0.03}$ $1.048^{*}_{\pm 0.18}$ $0.310_{\pm 0.05}$ $0.201^{*}_{\pm 0.01}$ $0.385_{\pm 0.04}$ $0.588_{\pm 0.05}$
Chain-of-Thought $0.089^{*}_{\pm 0.01}$ $0.145_{\pm 0.03}$ $0.881^{*}_{\pm 0.15}$ $0.340_{\pm 0.05}$ $0.203^{*}_{\pm 0.01}$ $0.385_{\pm 0.04}$ $0.578_{\pm 0.05}$
Socratic Prompting $084^{*}_{\pm 0.01} 0.0$$091_{\pm 0.03}$ $0.821^{*}_{\pm 0.14}$ $0.310_{\pm 0.05}$ $0.105^{*}_{\pm 0.01}$ $0.392_{\pm 0.04}$ $0.608_{\pm 0.05}$
SubQ Decomposition $0.085^{*}_{\pm 0.01}$ $0.118_{\pm 0.03}$ $1.058^{*}_{\pm 0.17}$ $0.320_{\pm 0.05}$ $0.190^{*}_{\pm 0.01}$ $0.405_{\pm 0.04}$ $0.549_{\pm 0.05}$
@@
Table A4: Model ablation study for evaluation pipeline. All experiments are run with vailla prompting. Across domains, model rankings are largely consistent across evaluators, indicating the robustness of our evaluation pipeline. Comparison of base models (explanation generator LLMs) across evaluator models (pipeline LLMs) for one dataset per domain. Best per evaluator is in bold; second-best is underlined.
Evaluation Model Gemini-2.5-Flash-Lite GPT-5-Mini Qwen2.5-VL-7B-Instruct
MassMaps
GPT-5.2-Pro $0.749\pm 0.011$ $0.901\pm 0.008$ $0.653\pm 0.020$
GPT-5-mini $0.644\pm 0.014$ $0.844\pm 0.011$ $0.570\pm 0.017$
Claude-Opus-4.5 $0.558\pm 0.018$ $0.774\pm 0.017$ $0.468\pm 0.016$
Claude-Haiku-4.5 $0.452\pm 0.018$ $0.742\pm 0.015$ $0.485\pm 0.017$
Gemini-2.5-Pro $0.793\pm 0.013$ $0.801\pm 0.012$ $0.629\pm 0.017$
Gemini-2.5-Flash $0.678\pm 0.014$ $0.836\pm 0.011$ $0.691\pm 0.017$
Cholecystectomy
GPT-5.2-Pro $0.681\pm 0.022$ $0.538\pm 0.030$ $0.686\pm 0.022$
GPT-5-mini $0.777\pm 0.015$ $0.757\pm 0.016$ $0.643\pm 0.019$
Claude-Opus-4.5 $0.665\pm 0.011$ $0.753\pm 0.011$ $0.666\pm 0.014$
Claude-Haiku-4.5 $0.625\pm 0.012$ $0.677\pm 0.010$ $0.594\pm 0.012$
Gemini-2.5-Pro $0.666\pm 0.012$ $0.697\pm 0.012$ $0.617\pm 0.014$
Gemini-2.5-Flash $0.660\pm 0.012$ $0.685\pm 0.012$ $0.594\pm 0.012$
Emotion
GPT-5.2-Pro $0.729\pm 0.019$ $0.906\pm 0.012$ $0.728\pm 0.028$
GPT-5-mini $0.730\pm 0.020$ $0.867\pm 0.019$ $0.751\pm 0.025$
Claude-Opus-4.5 $0.796\pm 0.019$ $0.916\pm 0.016$ $0.752\pm 0.027$
Claude-Haiku-4.5 $0.751\pm 0.024$ $0.881\pm 0.015$ $0.807\pm 0.022$
Gemini-2.5-Pro $0.806\pm 0.023$ $0.888\pm 0.018$ $0.723\pm 0.029$
Gemini-2.5-Flash $0.772\pm 0.026$ $0.900\pm 0.016$ $0.743\pm 0.028$
@@ Appendix C Ablation for Evaluation Models
To assess the robustness of our evaluation pipeline, we conduct ablations using different models as evaluators: closed-sourced GPT-5-mini (OpenAI, 2025b), Gemini-2.5-Flash-Lite (team, 2025), and an open-source model Qwen2.5-VL-7B-Instruct (Bai et al., 2025). We evaluate vanilla prompts on three datasets, each representing a different domain: Mass Maps (cosmology), Emotion (psychology), and Cholecystectomy (medicine).
Table˜A4 summarizes the results:
  • Mass Maps (cosmology): All three evaluators unanimously decide the explanations from two claude models, Claude-Opus-4.5 and Claude-Haiku-4.5, as the least aligned with expert intuition, while GPT and Gemini models align better.
  • Cholecystectomy (medicine): All three models give low $\mathsf{T\textrm{-}FIXscore}$s to Claude-Haiku-4.5, Gemini-2.5-Pro, and Gemini-2.5-Flash. All three evaluation models give highest and second highest scores among two of GPT-5.2-Pro, GPT-5-Mini, and Claude-Opus-4.5.
  • Emotion (psychology): All evaluators assign quite high scores across generation models, leading to minimal differences in rankings. On average, Claude-Opus-4.5 is consistently ranked as one of the top two models.
@@
Appendix D Expert Validation
D.1 Mass Maps
D.2 Supernova
D.3 Politeness
D.4 Emotion
D.5 Laparoscopic Cholecystectomy
Experts generally agreed with the model when it judged purely positional claims as not aligned with expert reasoning. They emphasized that referring only to a portion of the screen does not constitute surgically meaningful evidence for safety or risk. For example, the claim “The unsafe region occupies the lower to center-right portion of the image” was considered uninformative, as it describes only screen location rather than anatomically or surgically relevant cues. In such cases, experts agreed that alignment scores correctly reflected the lack of expert-valid reasoning.
Disagreement arose when the model assigned alignment to an incorrect anatomical category or relied on implausible visual inferences. For instance, experts judged the statement “The lower to center-right portion of the image shows darker, more vascular tissue” as less related to the “Peritoneal plane” category and more indicative of possible porta hepatis or vascular structures—although they emphasized that these structures cannot be directly identified from the image. Experts also noted that some unsafe-zone criteria, such as using the cystic lymph node as a landmark, are not visible in all cases and therefore difficult to apply consistently. In several instances, they pointed out that the model hallucinated tissues or structures that were not actually present in the image.
At a higher level, experts appreciated that the model correctly rejected claims based solely on screen position as misaligned with surgical intuition. However, they found the model less reliable when reasoning about anatomical structures whose visibility or relevance was uncertain. They stressed that many key anatomical landmarks cannot be directly observed in laparoscopic views and that some structures are far more informative than others for safety assessment. In particular, they highlighted the cystic lymph node and the estimation of the gallbladder infundibulum as useful landmarks, while cautioning against overinterpreting poorly visible or anatomically ambiguous regions.
D.6 Cardiac
D.7 Sepsis
Appendix E Prompts for T-FIX Pipeline
We show the prompts for Stage 1, 2, and 3 in Figure˜A1, Figure˜A2, Figure˜A3, and Figure˜A4, respectively. These prompts show a high-level template that was used by all domains. In practice, authors iterated multiple times on each domain’s prompts, experimenting with the instruction wording and few-shot examples that yielded the best possible results.
Appendix F Framework Development Process
In this section, we detail how we engineered all the prompts we tried for each of the 4 prompts in our T-FIX pipeline, what worked and what did not work, and how expert feedback was incorporated throughout the process.
F.1 Atomic claim extraction
The atomic claim extraction step decomposes each model-generated explanation into a set of atomic, independently interpretable claims for downstream relevance filtering and expert alignment. We designed structured prompts for each domain that enforce atomicity, standalone interpretability, and faithfulness to the original explanation.
What did not work. Given we began with the refined atomic claim extraction prompts from Wanner et al. (2024), this step needed little prompt engineering to work as expected.
What worked. We found that explicitly defining atomicity, standalone interpretation, and faithfulness within the prompt substantially improved the quality and consistency of extracted claims. Providing concrete in-context examples of correctly decomposed explanations for each domain further stabilized output structure and reduced stylistic variability across runs. Additionally, instructing the model to output only a list of claims with no additional commentary simplified downstream processing and facilitated automated relevance filtering and alignment.
F.2 Relevancy filtering
In the relevancy filtering step, we determine whether each extracted claim should be retained as supporting evidence for the model’s prediction. We prompted the LLM to label a claim as relevant only if it (i) is supported by the provided input and (ii) directly contributes to explaining the predicted outcome. This formulation filters out unsupported statements, generic background knowledge, and tautological claims that restate the prediction without adding evidence, yielding a cleaner set of claims for downstream grouping and expert alignment.
What did not work. Prompts that framed relevance in vague terms (e.g., “important,” “useful,” or “related”) led to inconsistent decisions and frequent retention of statements that were not meaningful for the explanation — restating the prediction, describing the input, describing the task, etc. We required multiple rounds of prompt iterations and a clear definition of relevance to get this stage to work with high accuracy and reliability.
What worked. We found that defining relevance with two explicit criteria, (i) the claim must be supported by the input and (ii) it must help explain the model’s prediction, yielded the most consistent filtering behavior. Requiring the model to generate a a reasoning to support its decision improved performance and made failure modes easier to diagnose during expert review. Providing a set of diverse examples further helped the model decide which claims were speculative, unsupported by the input, or not meaningful for the explanation.
F.3 Step 3: Claim grouping
This step required the highest number of iterations, as it is where we most incorporated domain expert feedback. We have broken down our process to the main components of prompt generation for expert knowledge.
Identifying Expert Categories. First, we used o3 deep research (OpenAI, 2025c) to seed our list of expert categories for each domain. We verified the model was pulling from reputable, existing literature, and manually ensured all categories made semantic sense and covered the areas that the experts expected. Following this, there were multiple (at least 2 per domain) iterations of review with the domain experts themselves for the determination of what constitutes the “ideal” expert categories for each domain. Working with the experts, we facilitated iterations where categories were modified, removed, and replaced, until experts reached consensus and the categories were finalized.
What did not work. Our initial setup of this pipeline did not involve the claim grouping step at all; we evaluated alignment of individual claims and aggregated the results. However, our domain experts pointed out that expert alignment is often determined by the joint content of multiple claims: claims can depend on one another, and some criteria are inherently compositional. For example, SIRS positivity in sepsis diagnosis requires meeting at least two symptoms from a predefined list, which may be expressed across several separate claims. Incorporating this feedback, we refactored our pipeline to incorporate claim grouping before expert alignment. This way, all claims related to a category could be judged together. Initial prompts returned groupings that were too selective; claims rarely belonged to multiple categories and many categories had zero related claims. Without explicit instructions to consider all claims before selecting a subset, GPT-5-mini sometimes defaulted to selecting only the most obviously relates claims and ignored supporting, but less obvious evidence.
What worked. After finalizing the expert criteria, we developed prompts that allowed us to group all relevant, filtered claims into criteria-level groups. We allowed the claims to belong to multiple criteria, as certain facts or statements in our explanations were referenced by multiple criteria. What worked well was curating a diverse set of examples with varying downstream groupings: most claims in a category, around 1/2 of the claims in a category, and no claims in a category. GPT-5-mini erred on the side of fewer claims in a category, so we included more examples of categories with large claim groupings to account for this. Upon encouraging GPT-5-mini to be more generous and inclusive with the grouping, we were satisfied with the results.
F.4 Expert alignment of claim groups
This step also required a high amount of iteration, and domain expert feedback was heavily used in creating our final alignment prompts.
What did not work. The first attempt of our claim to expert category alignment mapping was straightforward and simple. For each claim, we assigned it to the category that it most aligned with according to the LLM, and then calculated alignment. If the claim was completely not aligned with any of the expert categories, then it was omitted. We originally used a 0 to 1 scale with increments of 0.1 to score how aligned a claim was to a given expert category, where 1 is completely aligned and 0 is not aligned at all. This turned out to be too fine-grained a measure because it was unclear and ambiguous how to determine the difference between claims that got scores of 0.4 and 0.6, for example Moreover, numeric scores encouraged spurious precision and made it difficult for domain experts to interpret or validate the assignments. We also found that even with explicit definitions of the scale, the LLM’s scoring was unstable across runs and sensitive to minor prompt changes.
What worked. Based on feedback from our experts, we added the claim grouping step (see above). This group-based formulation better captured how expert reasoning often relies on multiple supporting statements rather than isolated claims. After experimentation, we also determined that using discrete alignment labels: none, partial, and complete alignments for each claim group was the most effective approach, because it provided clear semantic interpretation of how well a claim reflected an expert category. This formulation reduced ambiguity and improved consistency compared to numeric scoring schemes.
Examples of claim groups with partial alignment were especially important in helping GPT-5-mini handle borderline cases where claims were related to a category but lacked sufficient specificity. Similar to previous steps, requiring the LLM to provide a short justification for each alignment decision helped debug errors and improved performance.
What did not work. The initial versions of this prompt frequently resulted in the inclusion of too many claims per category, including irrelevant or weakly related statements. The LLM also tended to omit subtle but important claims when they were phrased indirectly or lacked explicit category keywords. Additionally, without explicit instructions to consider all claims before selecting a subset, the model sometimes defaulted to selecting only the most obvious claims and ignored secondary evidence. This required prompt revisions that enforced exhaustive consideration of the claim set and stricter inclusion criteria.
@@ F.4 Expert alignment of claim groups
Figure A2: Prompt Template for Stage 2A: Relevancy Filtering
Prompt
Figure A3: Prompt Template for Stage 2B: Claim Grouping
@@
Appendix G T-FIX Datasets: Additional Details
@@ G.1 Mass Maps
Data Selection & Preprocessing. We randomly sampled 100 examples from the MassMaps test set. To ensure compatibility with LLMs like GPT-5-mini, which operate on a 32×32 patch size, we upsampled each image by a factor of 11 to preserve spatial detail and avoid patch-level compression. Instead of raw pixel values, we applied a colormap based on expert-defined intensity thresholds used to identify key cosmological features such as voids and clusters. Pixel intensities were scaled by standard deviations to emphasize meaningful variation. We found that larger, visually enhanced inputs reduced refusal rates from LLMs and encouraged more consistent responses.
Explanation Prompt. Figure A7 shows the prompt used to generate LLM explanations for predicting $\Omega_{m}$ and $\sigma_{8}$. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6. The prompt includes a description of how pixel values are mapped to colors, as well as the valid ranges for $\Omega_{m}$ and $\sigma_{8}$. Without this range, models tend to default to common values (e.g., 0.3 for $\Omega_{m}$, 0.8 for $\sigma_{8}$), reducing response variability.
Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below:
@@
G.2 Supernova
Task. The objective is to classify astrophysical objects using time-series data comprising observation times (Modified Julian Dates), wavelengths (filters), flux values, and corresponding flux uncertainties. We use data from the PLAsTiCC challenge (Team et al., 2018), where the model must predict one of 14 astrophysical classes.
Data Selection & Preprocessing. We sampled 100 examples across the Supernova train, validation, and test sets, aiming for 7–8 instances per class to mitigate class imbalance. For rare classes with only one test set instance, we included all available examples from the validation and test sets, supplementing with training samples to meet the target count. For LLM input, we converted each raw time series into a multivariate time-series plot: time is on the x-axis, flux on the y-axis, error bars denote flux uncertainty, and point colors indicate different wavelengths.
Explanation Prompt. Figure A8 shows the prompt used to generate explanations for classifying astronomical objects. We replace [BASELINE_PROMPT] with one of four prompting strategies shown in Figure˜A6. The prompt includes a description of the input plot as a multivariate time series and provides the full list of possible class labels to guide the model’s predictions.
Expert Criteria. The expert-validated criteria for expert alignment calculation are listed below: